EVN experiment feedback facility: N20C1


Session: Feb 2020 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Irbene (Roberts at Tue Feb 25 15:08:26 2020) Success:

Hartebeesthoek (Jonathan Quick at Tue Feb 25 15:37:52 2020) Success:
No known problems.

Svetloe (Isaenko O. at Tue Feb 25 17:15:15 2020) Success:

Badary (Alakova, Plotnikova at Tue Feb 25 17:15:30 2020) Success:

Zelenchukskaya (Lysenkova M. at Tue Feb 25 17:15:59 2020) Success:

Onsala (Jun Yang at Wed Feb 26 09:48:08 2020) Success:

Yebes (Daniel Sastre Walter at Wed Feb 26 14:04:50 2020) Success:
Not known problems.

Effelsberg (Uwe at Wed Feb 26 14:50:15 2020) Minor failures: HIGH
When switching the source in the middle of the observations the DBBC lost somehow the calibration on IFC. All LCP data shows an enormous increase in Tsys and huge spikes: e.g. http://old.evlbi.org/tog/ftp_fringes/N20C1/scan08/st04_lcp-st04_lcp-ch0-lsb-0-2_large.png Reason not understood yet.

Tianma65 (Bo Xia at Thu Feb 27 05:04:32 2020) Failure:
didn't observe due to the problem with fibre transmission unit.

Torun (Pawel Wolak at Fri Feb 28 11:20:26 2020) Minor failures:
No cal signal.

Westerbork (Blaauw at Fri Mar 06 10:34:52 2020) Success: